We wish to check that normal body temperature may be less than 98.6 degrees. In a random sample of n = 18 individuals, the sample mean was found to be 98.217 and the standard deviation
was 0.684. Assume the population is normally distributed. Use=0.05.
1
Expert's answer
2020-12-02T18:15:57-0500

Population standard deviation is unknown.

"H_0:\mu=98.6"

"H_a:\mu<98.6"

"t=\frac{\bar X-\mu}{\frac{s}{\sqrt{n}}}"

"t=\frac{98.217-98.6}{\frac{0.684}{\sqrt{18}}}"

"=-2.376"

"t_{0.05,17}=-1.74"

Since the test statistic -2.376 is less than the critical value -1.74, we reject the null hypothesis and conclude that there is enough evidence at 95% confidence level to conclude that normal body temperature is less than 98.6 degrees.